I was asked to describe three rules for exponents and came up with the following:
Product Rule – When multiplying numbers with the same base we add the exponents together
Quotient Rule – When dividing numbers with the same base we keep the base, subtracting the exponent of the denominator from the exponent of the numerator
Power Rule – This is where we raise the base and multiply the exponents
Will those descriptions suffice? Also can you please provide me with an example of an expression that uses scientific notation and at least one of the rules for exponents that I have described? Please and thank you.

Answer by stanbon(75887)

I was asked to describe three rules for exponents and came up with the following:
Product Rule – When multiplying numbers with the same base we add the exponents together.
Good
-----------------------
Quotient Rule – When dividing numbers with the same base we keep the base, subtracting the exponent of the denominator from the exponent of the numerator
I would say: "Subtract the smaller exponent from the larger exponent".
-----------------------
Power Rule – This is where we raise the base and multiply the exponents
I would say: We raise the base to a power equal to the product of the
exponents.
------------------------
Also can you please provide me with an example of an expression that uses scientific notation and at least one of the rules for exponents that I have described?
(3.2x10^2)/(1.6x10^5) = (3.2/1.6)(10^2/10^5) = 2*10^(-3) = 0.002
=================================
Cheers,
Stan H.